OpenStudy (anonymous):

The variables x, v, and a have the dimensions of [L], [L]/[T], and [L]/[T]^2, respectively. These variables are related by an equation where n is an integer constant (1, 2, 3, etc.) without dimensions. What must be the value of n, so that both sides of the equation have the same dimensions? Explain your reasoning.

OpenStudy (anonymous):

The equation is v^n = 2ax

OpenStudy (anonymous):

I got to:\[\frac{ L }{ T }^{n}=2\frac{ L }{ T }^{a}\times L\]

OpenStudy (anonymous):

=L^2/T^2 So n has to equal 2 right?

OpenStudy (kirbykirby):

yes I get n=2